# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ 3.3.1 Color sum profiling
Initially, I profiled the code by a sampling approach, using perf, but this did not allow detailed color sum profiling on the GPU. I therefore made some additional modifications to the ihel3 version of the code to instrument it with dedicated timers. Specifically, I used some timers based on the x86 rdtsc instruction, which I had developed for some previous profiling work on MG5aMC bib:prof2024 . This approach can provide relatively accurate results with a limited (<<10%) overhead.
The results of my analysis, which are shown in Fig. 7 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 6% and 15% in double, mixed and single precision modes, respectively. These results are particularly good for the mixed precision mode, and indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation.
In Fig. 7, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. These results are instead shown by the two red curves in Fig. 8, which show that the efficiency of BLAS color sums increases as nevt increases, while that of the kernel implementation is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it. The two other black curves in Fig. 8 refer to the further improvements in the “ihel3p1” version of the software, described further below.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 after some discussions in the team (and notably some suggestions by Olivier Mattelaer), and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation, with a factor 2. While writing this paper, I realized that this optimization had been applied to C++ code, but not yet to color sum kernels on GPUs: this is another point that I later fixed in the “ihel3p1” version, described further below.
@@
Table 3: Summary of the results for the profiling of the check.exe application, using the NVidia NSight Compute (ncu) profiler version 2024.1.1.0. The results are obtained on an NVidia A100 GPU, using mixed floating precision in the ihel3 version of CUDACPP. Only the kernels in one specific helicity stream were profiled. The columns represent the following: (1) color sum implementation (three configurations are tested: default using CUDA kernels; cuBLAS; cuBLAS with TF32 math mode); (2) GPU grid size configuration for the kernel implementation (or, equivalently, number of events processed in one cuBLAS multiplication); (3) simplified function name printed out by the profiler; (4) FMA (CUDA core) activity, as per ncu metric sm__pipe_fma_cycles_active.avg.pct_of_peak_sustained_active; (5) Tensor core activity, as per ncu metric sm__pipe_tensor_cycles_active.avg.pct_of_peak_sustained_active.
@@ 3.3.2 Tensor core profiling
As the use of tensor cores was one of the main motivations for prototyping cuBLAS for color sums, I present in Table 3 the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the results shown so far in this paper, these were obtained using an NVidia A100 GPU, which has more advanced tensor core features than the older V100 model. The results of this test, which was performed using CUDACPP ihel3 with mixed floating precision, clearly show that tensor cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of tensor cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of tensor cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of tensor cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as tensor cores are designed and optimized for lower precision calculations in AI. A more detailed investigation of tensor cores in the CUDACPP cuBLAS implementation is beyond the scope of this paper, but could be useful for instance in the context of the HEP-SCORE benchmarking project. In particular, further investigations could be useful to understand the usefulness of cuBLAS for CUDACPP color sums on more recent NVidia GPUs, with much more advanced tensor core features.
3.4 Further improvements of color sums (“ihel3p1”)
@@
As mentioned above, while writing the first preprint for this paper bib:preprintV1 I realised that two optimizations of color sums on GPUs could still be attempted. I have implemented both of them in a later version “ihel3p1” of the code, the results for which are represented by the black-solid (ihel3p1, BLAS disabled) and black-dashed lines (ihel3p1b, BLAS enabled) in Figures 4 and 5 for NVidia and AMD GPUs. These patches have already been merged in production with the rest of my first batch of changes to the code: the black lines in the figures, in other words, represent the performance of the current latest release CUDACPP v1.01.01. As shown in Fig. 6, I also note that the C++ implementation in this ihel3p1 release has essentially the same performance as the previous ihel0 release for all SIMD builds.
@@
The second optimization, conversely, concerns BLAS and was motivated by the observations, discussed above and based on Figures 4 and 8, that BLAS throughputs significantly increase when large vectors are used. The basic idea, therefore, was to group together the color sums for all helicities as a single BLAS multiplication, rather than perform a separate BLAS multiplication for each helicity. The changes that I implemented are therfore the following: first, I removed the color sum dispatcher from the helicity loop, implementing a single dispatcher for all helicities rather than one per helicity; in the default case where color sum kernels are used, the dispatcher takes care of launching separate kernels for the different helicities, in the corresponding helicity streams; if BLAS is selected at runtime, however, the dual amplitudes from the different helicities are concatenated to form a single much larger array of events, as if the different helicities for each event were treated as different events, and a single BLAS multiplication is performed in the default stream. This is represented in the lower diagrams of Fig. 11 (this refers to a later ihel6 version of the software, which however only differs from ihel3p1 in the handling of Feynman diagrams). As shown in Fig. 4, the BLAS throughputs of this improved ihel3p1b implementation (dashed-black) are much higher than those of the previous ihel3b implementation (dashed-red). For $gg\!\rightarrow\!t\bar{t}ggg$ (bottom row), in particular, the performance receives a massive boost: BLAS color sums are now better than kernels in double precision, and have equivalent performance in mixed and single precision. This can also be seen in Fig. 8, which shows that the fraction of time spent in the color sum in $gg\!\rightarrow\!t\bar{t}ggg$ is now consistently in the 5–10% range when BLAS is used, independently of the number of events in each GPU iteration. For simpler processes, however, color sum kernels still maintain a higher performance: keeping this into account, it seems appropriate to keep color sum kernels as the default and BLAS as an optional alternative.

### v3

@@ 3.3.1 Color sum profiling
Initially, I profiled the code by a sampling approach, using perf, but this did not allow detailed color sum profiling on the GPU. I therefore made some additional modifications to the ihel3 version of the code to instrument it with dedicated timers. Specifically, I used some timers based on the x86_64 (Read Time-Stamp Counter) rdtsc instruction, which I had developed for some previous profiling work on MG5aMC bib:prof2024 . This approach can provide relatively accurate results with a limited (<<10%) overhead.
The results of my analysis, which are shown in Fig. 7 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 6% and 15% in double, mixed and single precision modes, respectively. These results are particularly good for the mixed precision mode, which may indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation.
In Fig. 7, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. These results are instead shown by the two red curves in Fig. 8, which show that in ihel3 the efficiency of BLAS color sums improves as nevt increases, while for the kernel implementation it is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it. The two other black curves in Fig. 8 refer to the further improvements in the “ihel3p1” version of the software, described below.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 after some discussions in the team (and notably some suggestions by Olivier Mattelaer), and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation, with a factor 2. While writing this paper, I realized that the latter optimization had been applied to color sums in C++, but not yet in GPU kernels: this is another point that I later fixed in the “ihel3p1” version, described further below.
@@
Table 3: Summary of the results for the profiling of the check.exe application, using the NVidia NSight Compute (ncu) profiler version 2024.1.1.0. The results are obtained on an NVidia A100 GPU at CERN, on a node equipped with AMD EPYC 7313 CPUs, using mixed floating precision (FPTYPE=m) in the ihel3 version of CUDACPP. Code built using CUDA 12.4 and gcc 11.5. Only the kernels in one specific helicity stream were profiled. The columns represent the following: (1) color sum implementation (three configurations are tested: default using CUDA kernels; cuBLAS; cuBLAS with TF32 math mode); (2) GPU grid size configuration for the kernel implementation (or, equivalently, number of events processed in one cuBLAS multiplication); (3) simplified function name printed out by the profiler; (4) FMA (CUDA core) activity, as per ncu metric sm__pipe_fma_cycles_active.avg.pct_of_peak_sustained_active; (5) tensor core activity, as per ncu metric sm__pipe_tensor_cycles_active.avg.pct_of_peak_sustained_active.
@@ 3.3.2 Tensor core profiling
As the use of tensor cores was one of the main motivations for prototyping cuBLAS for color sums, I present in Table 3 the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the results shown so far in this paper, these were obtained using an NVidia A100 GPU, which has more advanced tensor core features than the older V100 model. This test, performed using CUDACPP ihel3 with mixed floating precision, clearly shows that tensor cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of tensor cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of tensor cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of tensor cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as tensor cores are designed and optimized for lower precision calculations in AI. A more detailed investigation of tensor cores in the CUDACPP cuBLAS implementation is beyond the scope of this paper, but could be useful for instance in the context of the HEP-SCORE benchmarking project. Further investigations of cuBLAS performance for CUDACPP color sums on more recent NVidia GPUs, with much more advanced tensor core features, could also be useful.
$gg\!\rightarrow\!t\bar{t}ggg$ — current code (master/ihel3p1)
C++ SIMD build mode none novec-cs none sse4 avx2 512y 512z
FPTYPE=d Time to compute 128 MEs 1.293s 0.680s 0.305s 0.268s 0.152s
$\rightarrow$ Time in calculate_jamps 1.233s 0.626s 0.280s 0.243s 0.138s
$\rightarrow$ Time in color_sum_cpu 0.061s 0.054s 0.025s 0.025s 0.013s
FPTYPE=m Time to compute 128 MEs 1.308s 0.646s 0.289s 0.252s 0.143s
$\rightarrow$ Time in calculate_jamps 1.232s 0.616s 0.275s 0.238s 0.136s
$\rightarrow$ Time in color_sum_cpu 0.076s 0.031s 0.014s 0.014s 0.008s
FPTYPE=f Time to compute 128 MEs 1.242s 0.304s 0.152s 0.133s 0.075s
$\rightarrow$ Time in calculate_jamps 1.206s 0.276s 0.139s 0.121s 0.069s
$\rightarrow$ Time in color_sum_cpu 0.037s 0.028s 0.012s 0.013s 0.006s
$gg\!\rightarrow\!t\bar{t}ggg$ — new code (csm)
C++ SIMD build mode none novec-cs none sse4 avx2 512y 512z
FPTYPE=d Time to compute 128 MEs 1.336s 1.286s 0.680s 0.305s 0.267s 0.152s
$\rightarrow$ Time in calculate_jamps 1.229s 1.230s 0.626s 0.280s 0.242s 0.139s
$\rightarrow$ Time in color_sum_cpu 0.108s 0.056s 0.054s 0.025s 0.025s 0.013s
FPTYPE=m Time to compute 128 MEs 1.340s 1.264s 0.647s 0.288s 0.251s 0.144s
$\rightarrow$ Time in calculate_jamps 1.232s 1.230s 0.617s 0.275s 0.238s 0.137s
$\rightarrow$ Time in color_sum_cpu 0.107s 0.034s 0.029s 0.014s 0.014s 0.007s
FPTYPE=f Time to compute 128 MEs 1.313s 1.240s 0.305s 0.152s 0.133s 0.076s
$\rightarrow$ Time in calculate_jamps 1.206s 1.207s 0.276s 0.139s 0.121s 0.069s
$\rightarrow$ Time in color_sum_cpu 0.107s 0.033s 0.028s 0.013s 0.013s 0.006s
Table 4: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}ggg$ on CPU for different SIMD modes and floating point precisions, in the current version (top: ihel3p1) and an improved version (bottom: csm) of CUDACPP. The code is instrumented with rdtsc timers for color sum profiling. Results for a single core of an Intel Xeon Gold 6326 CPU at CERN (using gcc 11.5).
3.4 GPU color sum further improvements (“ihel3p1”)
@@
As mentioned above, while writing the first preprint for this paper bib:preprintV1 I realised that two optimizations of color sums on GPUs could still be attempted. I have implemented both of them in a later version “ihel3p1” of the code, the results for which are represented by the black-solid (ihel3p1, BLAS disabled) and black-dashed lines (ihel3p1b, BLAS enabled) in Figures 4 and 5 for NVidia and AMD GPUs. These patches have already been merged in production with the rest of my first batch of code changes: the black lines in the figures, in other words, represent the performance of the current latest release CUDACPP v1.01.01. As shown in Fig. 6, I also note that the C++ implementation in this ihel3p1 release has essentially the same performance as the previous ihel0 release for all SIMD builds.
@@
The second optimization, conversely, is specific to BLAS and was motivated by the previous observations, based on Figures 4 and 8, that BLAS throughputs significantly increase when large vectors are used. The basic idea, therefore, was to group together the color sums for all helicities as a single BLAS multiplication, rather than perform a separate BLAS multiplication for each helicity. The changes that I implemented are therfore the following: first, I removed the color sum dispatcher from the helicity loop, implementing a single dispatcher for all helicities rather than one per helicity; in the default case where color sum kernels are used, the dispatcher takes care of launching separate kernels for the different helicities, in the corresponding helicity streams; if BLAS is selected at runtime, however, the dual amplitudes from the different helicities are concatenated to form a single much larger array of events, as if the different helicities for each event were treated as different events, and a single BLAS multiplication is performed in the default stream. This is represented in the lower diagrams of Fig. 11 (this refers to a later ihel6 version of the software, which however only differs from ihel3p1 in the handling of Feynman diagrams). As shown in Fig. 4, the BLAS throughputs of this improved ihel3p1b implementation (dashed-black) are much higher than those of the previous ihel3b implementation (dashed-red). For $gg\!\rightarrow\!t\bar{t}ggg$ (bottom row), in particular, the performance receives a massive boost: BLAS color sums are now better than kernels in double precision, and have equivalent performance in mixed and single precision. This can also be seen in Fig. 8, which shows that the fraction of time spent in the color sum in $gg\!\rightarrow\!t\bar{t}ggg$ is now consistently in the 5–10% range when BLAS is used, independently of the number of events in each GPU iteration. For simpler processes, however, color sum kernels still have a higher performance: taking this into account, it seems appropriate to keep color sum kernels as the default and BLAS as an optional alternative.
3.5 CPU color sum SIMD optimizations (“csm”)
In this section, I describe some recent studies and optimizations of color sums specifically targeting their implementation on vector CPUs. This work, which corresponds to the orange box in Fig. 1, represents my most recent and final addition to this paper, and was only included after my second preprint bib:preprintV2 of it. Its main motivation, in fact, was a desire to better understand the SIMD speedups of $gg\!\rightarrow\!t\bar{t}gggg$ that I initially reported in Ref. bib:preprintV2 and which are listed later on in Table 7 in Sec. 4.4.1. In particular, in mixed precision mode, where helicity amplitudes are computed in double precision and color sums in single precision, the SIMD speedups of both calculations with respect to the none build mode seem consistent with that expected for double precision (e.g. a factor 2 from none to sse4), while a larger speedup (a factor 4 for sse4) would be expected for color sums. Based on previous observations, I initially suspected that this may be due to a suboptimal SIMD implementation of FPTYPE=m color sums, which led me to investigate them in more detail. Eventually, I found that mixed-precision color sums on CPUs already have a reasonable performance, except in some corner cases which I improved. I understood, instead, that the reason for the lower-than-expected SIMD speedups with respect to the none mode is that the latter does use some SIMD through auto-vectorization, and cannot always be taken as a reference for no-SIMD performance. I give more details in the following.
To start with, I analysed the absolute time spent in the Feynman diagram and color sum calculations in $gg\!\rightarrow\!t\bar{t}ggg$ for different SIMD build modes and floating point precisions, in the current version of CUDACPP (ihel3p1). The results, which are listed in the top half of Table 4, are consistent with and complementary to those previously given in Fig. 7, where only the fraction of time taken by color sums was shown. The main takeaway is that, for all SIMD modes (sse4 to 512z), the time spent in color sums in mixed precision is only 10% to 20% larger than in single precision, and almost a factor 2 smaller than in double precision. This is an indication that the SIMD implementation of FPTYPE=m color sums does have a reasonably good performance, and there is only limited room to further optimize it to bring it closer to that of FPTYPE=f. Two other observations concern the none build mode: to start with, none color sums are twice slower in FPTYPE=f with respect to FPTYPE=f, while they should normally be just as fast; in addition, FPTYPE=f color sums are only 30% slower for none with respect to sse4, while one would expect them to be a factor 4 slower. It should be noted, in particular, that the last two observations only concern color sums: in fact, the Feynman diagram calculations in none mode all take the same time for FPTYPE=d,m,f, and are around a factor 2,2,4 slower than sse4, respectively, as expected. Based on these observations, I performed three further studies and optimizations in a new “csm” branch of the code, as described below.
A first optimization consisted in streamlining the way the actual multiplication $V^{t}(C)V$ is performed for a given precision (i.e. using double precision vectors $V$ and matrices $C$ for FPTYPE=d and single precision $V$ and $C$ for for FPTYPE=m,f). As shown in the bottom half of Table 4, this change speeds up all build modes, and especially the none build mode, where FPTYPE=m color sums are twice faster than they were and are now as fast as for FPTYPE=f.
A second test consisted in explicitly disabling auto-vectorization in the CPU color sum, using #pragma GCC optimize("no-tree-vectorize"). The results of this test, shown in Table 4, clearly show that color sums are significantly slower in this modified “novec-csnone mode than in the default none build mode, and are now around a factor 2,2,4 slower than sse4, respectively, for FPTYPE=d,m,f. This is an indication that the -march=x86-64 flag in the none mode does not completely disable SIMD operations: in fact, the gcc manual bib:gccx86 states that -march=x86-64 specifies “a generic CPU with 64-bit extensions, MMX, SSE, SSE2, and FXSR instruction set support”, which does allow SSE and SSE2 SIMD instructions on 128-bit xmm vector registers. In retrospective, I recognize that it may have been a mistake to base the definition of the none build mode on -march=x86-64 alone, and it might have been better to also disable auto-vectorization, if the only purpose was that of having a no-SIMD reference scenario. On the other hand, the main point of the none build mode is that of testing software performance without any explicit event-level vectorization, i.e. when all operations are performed using scalar fptype variables rather than vectors fptype_v of neppV variables. It is also very interesting that a significant level of auto-vectorization in the none build mode only seems to happen in color sums, not in Feynman diagram calculations. This is an indication that the auto-vectorization of color sums in the none mode exploits color-level SIMD parallelism, as it is the multiplication of the color matrix and dual color amplitudes for a single event that gets auto-vectorized. This is different from what happens in the sse4, avx2, 512y and 512z modes, where the speedup comes — by design — from event-level SIMD parallelism.
Finally, a third set of tests and attempted optimizations I performed, which is very specific to FPTYPE=m color sums on CPUs, concerns not the actual calculation $V^{t}(C)V$ of the color sum in single precision, but the mechanisms to interface the double-precision outputs Feynman diagram calculation as the single-precision inputs to the color sum, and the single precision outputs of the color sum as the double-precision inputs to the further processing of matrix elements. As I briefly hinted in a footnote of Sec. 2.2, the algorithm that I implemented for the calculation of mixed-precision SIMD color sums is the following: the Feynman diagram calculation is performed sequentially for two vectors of neppV events, producing two fptype_v SIMD vectors of neppV double precision color amplitudes (jamp); an ad-hoc function fpvmerge converts these two fptype_v vectors into one SIMD vector fptype2_v of 2*neppV single precision color amplitudes; the color sum is performed in single precision, using event-level SIMD parallelism; its output is one SIMD vector fptype2_v of 2*neppV single precision squared matrix elements; two functions fpvsplit0 and fpvsplit1 extract the lower and higher halves of this vector as two fptype_v SIMD vectors of neppV double precision squared matrix elements. The current ihel3p1 version of CUDACPP still uses my original implementation of the fpvmerge and fpvsplit0/1 functions, which is based on CVE constructors from hardcoded initializer lists. Over time, I often suspected that this implementation was suboptimal and may cause a large overhead of FPTYPE=m color sums over FPTYPE=f. My third optimization in the csm branch simply consists in two new implementations of the fpvmerge function, one using Intel intrinsics bib:intrinsics and another using the std::experimental::simd proposed C++ extension. The results of some tests of these new implementations in $gg\!\rightarrow\!t\bar{t}ggg$, which I do not list in detail, are actually quite reassuring: my original implementation of fpvmerge is just as fast, within 1 or 2%, as that using intrinsics, and both are up to 30% faster than that using experimental SIMD.
In summary, this work was very useful to better understand the SIMD performance of CPU color sums, clarifying some issues that had previously appeared in $gg\!\rightarrow\!t\bar{t}gggg$ tests, and it also resulted in a few optimizations of SIMD color sums. The new code is available in a github tag bib:csmtag of the csm branch, from which I created a pull request that I recommend to merge upstream. In this pull request, the original implementation of the fpvmerge function based on initializer lists is kept as the default; the new versions based on intrinsics and experimental SIMD are disabled but are still available in the code, where they may provide a useful reference for future developments.
4 Kernel splitting II: Feynman diagram groups
After implementing helicity-level parallelism, separating Feynman diagrams from color sums and porting the latter to BLAS, the next step in my kernel splitting developments consisted in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. My initial motivation for attempting this was to investigate whether this may lead to further throughput increases, possibly thanks to the lower number of registers used by smaller GPU kernels (lower “register pressure”). Unfortunately, this strategy did not allow me to achieve further throughput increases; this is probably because moderately small Feynman diagram kernels are still quite complex and especially come at the cost of a large increase in GPU global memory access, which massively slows down the calculation. However, there is also a second motivation for splitting Feynman diagrams, namely making the code easier to build and manage for complex physics processes with very large numbers of diagrams. As described in this section, my work in this direction did achieve some interesting progress, such as the possibility to calculate MEs for $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes that were previously non-computable in the CUDACPP software.
In the following subsections, I describe my second batch of three sets of kernel splitting changes, which I now also consider ready to be integrated in new releases of CUDACPP: (ihel4) one kernel per Feynman diagram in each helicity stream; (ihel4p1) ditto, with optional CUDA graph orchestration; (ihel5 — and ihel6 with DCDIAG=0) one kernel per diagram group in each helicity stream, with separate source code files; (ihel6 with DCDIAG=1) a single kernel in each helicity stream, processing all diagram groups as device functions. The performance of the intermediate ihel4 step and of the final ihel6p2 step, for $gg\!\rightarrow\!t\bar{t}$ production with 0, 1, 2 or 3 extra final state gluons, are given in Figures 4 and 5 for NVidia and AMD GPUs. These plots show that the single-diagram kernel strategy of ihel4 is very inefficient, but that throughputs that are completely equivalent to those of the earlier ihel3p1 software are recovered in the newer ihel6p2 if the code is generated without diagram splitting, i.e. with a single kernel processing all Feynman diagrams as in ihel3p1. In other words, the usefulness of ihel6p2 is not that it yields higher throughputs for our standard candle processes; instead, its main interest is that it allows to compute the more complex $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ with 4 and 5 extra final state gluons, the results for which are presented at the end of this section.
4.1 Feynman diagrams as individual kernels (“ihel4”)